OpenStudy (anonymous):

What value of w solves the equation? w^-3 =1/64

OpenStudy (anonymous):

do you understand negative exponents?

OpenStudy (anonymous):

Yea.

OpenStudy (anonymous):

do you know your powers?

OpenStudy (anonymous):

Like the basic rules are that you have to simplify

OpenStudy (anonymous):

\[\begin{matrix}1^3=1 \\2^3=8\\ 3^3=27\\ 4^3 = ?\end{matrix}\]

OpenStudy (anonymous):

16? 4x4 is 16 just like the expont 4^2

OpenStudy (anonymous):

*exponet

OpenStudy (anonymous):

no, \(4^2 = 16\) what is \(4^3 = ?\)

OpenStudy (anonymous):

in decmials its 0.0625

OpenStudy (anonymous):

\(\large{4^3 =\, ?}\)

OpenStudy (anonymous):

0.015625

OpenStudy (anonymous):

64

OpenStudy (anonymous):